# HP-BERT: A framework for longitudinal study of Hinduphobia on social media via LLMs

### v1

@@ 3.1 Datasets
3.1.1 Hinduphobic COVID-19 X (Twitter) Dataset
We extracted the Hinduphobic tweets and curated a dataset using the X (Twitter) application programmer interface (API) [^2] https://developer.x.com/en/docs/x-api. The dataset spans from April 2020 to January 2024, capturing the context of the COVID-19 pandemic with tweets originating from India, featuring both Hinglish (a mix of Hindi and English) and English. Each tweet includes terms such as COVID, COVID-19, or coronavirus, establishing a clear relation between the pandemic and the Hindu community. We employed targeted keywords and hashtags to identify negative (Hinduphobic) and positive/neutral sentiments associated with Hinduism. We present some of the terms identified and their relevance to Hinduphobic sentiments in (Table 1) along with positive/neutral terms. The negative terms such as ”Hindu superspreaders” (referring to the Kumbh Mela Hindu festival), ”Hindu rituals causing COVID” (related to temple gatherings), and derogatory references such as ”piss drinker” (associated with cow urine and dung) were used to malign and blame the Hindu community for the spread of COVID-19. Notable hashtags promoting Hinduphobic content included #CoronaJihad, #HindusSpreadCorona, #HinduCOVIDConspiracy, #COVID19HinduHate, and #COVIDHindutva. For the positive or neutral sentiment, we focused on terms associated with Hindu contributions to COVID-19 relief efforts, such as ”Namaste” or ”Namaskar” (a Hindu greeting that avoids physical contact) and references to organisations such as ISKCON (International Society of Krishna Consciousness), RSS (Rashtriya Swayamsevak Sangh), Swaminarayan Temple, and Mahavir Temple. These organisations played significant roles in COVID-19 relief by providing food, plasma donations, hospital support, and financial aid. Positive hashtags in this category included #NamasteForSafety, #HindusForHumanity, #ISKCONRelief, #RSSHelpingHands, #TempleAidCOVID, and #HinduCOVIDRelief. We derived the full list of keywords from scholarly sources [96], Hindu American Foundation [^3] https://www.hinduamerican.org/hindu-hate-glossary, news articles [^4] https://shorturl.at/3Dvrl, and Wikipedia [^5] https://en.wikipedia.org/wiki/Anti-Hindu_sentiment with explanations for each term’s association with negative or positive sentiments available in a document [^6] https://github.com/pinglainstitute/Hinduphobia-COVID-19-beyond.
@@ 3.1.1 Hinduphobic COVID-19 X (Twitter) Dataset
Table 1: Keywords and Their Association with Negative (Hinduphobic) and Positive & Neutral (Pro-Hindu) Sentiments in Comments
@@
Hindus are the real reason COVID is spreading! Can’t believe they still gather at temples! #HinduSuperspreaders #COVID19 #Pandemic hindus are the real reason covid is spreading cannot believe they still gather at temples hindusuperspreaders covid19 pandemic
@@
The thinly veiled Hinduphobia peeping from behind ‘secularism’ facade is tiring and oh so cliche. #Secularism #Hinduphobia the thinly veiled hinduphobia peeping from behind secularism facade is tiring and oh so cliche secularism hinduphobia
@@
Table 2: Examples of tweets related to Hinduphobic before and after preprocessing.
S.No Original Preprocessed S.No Original Preprocessed S.No Original Preprocessed
1 :) happy 7 i’ll’ve I will have 13 @username [user mention]
2 :( sad 8 omg oh my god 14 plz please
3 ain’t am not 9 #COVID19 COVID-19 15 u you
4 lol laugh out loud 10 u2 you too 16 thx thanks
5 smile smile 11 rt retweet 17 idk I don’t know
6 #StaySafe StaySafe 12 asap as soon as possible 18 ik I know
Table 3: Transforming tweet with emojis to standardised language for data processing by language models.
Figure 2: BERT (HP-BERT) Model Architecture for Detecting Hinduphobic Sentiment, Featuring Tokenization (Encoding) of Hinglish/English Tweets Using a Set of Encoders. It Includes a Fully Connected Layer for Classifying Tweets as Hinduphobic or Non-Hinduphobic/Neutral, and the Final Layer Provides Sentiment Classification into Different Categories.
@@
After collecting and curating the Hinduphobic X (Twitter) dataset, we preprocessed it to enhance its quality and relevance by removing irrelevant content, spam, and duplicate entries, along with eliminating web links, hashtags, and profile tags [97, 98, 99]. We also expanded contractions and abbreviations to ensure consistency in data preparation. Additionally, we transformed the emojis into text along with other non-text symbols [100, 101] as shown in (Table 3). Furthermore, (Table 2) presents examples of tweets before and after preprocessing.
@@
We employed a Human-in-the-Loop (HITL) approach [102, 95] to ensure accurate labelling of the processed tweets. We manually labelled 2,000 tweets as either (1) negative sentiment (Hinduphobic) or (2) positive/neutral sentiment (pro-Hindu). These labelled examples were used as training data for the GPT-3.5-Turbo model, accessed via the OpenAI API [^7] https://platform.openai.com/docs/models/gpt-3-5-turbo. After GPT-3.5-Turbo labelled the remaining tweets, the automatically labelled tweets were manually checked. The final labelled dataset, containing 8,000 tweets (4,600 labelled as Hinduphobic and 3,400 as positive/neutral), was published on Kaggle [103]. The need for manual labelling stemmed from the challenge of accurately identifying hate speech, which is language and culture-dependent. Existing models, such as HateBERT [65] and COVID-HateBERT [104], may struggle to detect nuanced expressions of Hinduphobic content with hidden meanings. Correctly identifying instances of hate speech is essential to maintaining data integrity, ensuring accurate interpretation, and minimising the risk of inaccuracies. Figure 1 illustrates the process of dataset creation and processing. This dataset will be used to fine-tune the BERT (HP-BERT) model, a specialised model designed to improve the detection of Hinduphobic content.
3.1.2 SenWave dataset
@@
The SenWave dataset [79], contains over 104 million tweets and Weibo messages related to COVID-19 in six different languages: English, Spanish, French, Arabic, Italian, and Chinese. The data was collected between March 1 and May 15, 2020, to analyse social media posts surrounding the pandemic. It includes 10,000 English and 10,000 Arabic tweets labelled into ten sentiment categories: optimistic, thankful, empathetic, pessimistic, anxious, sad, annoyed, denial, official report, and joking. This comprehensive labelling facilitates in-depth analysis of the diverse emotional responses exhibited during the pandemic. The effectiveness of the SenWave dataset has been validated in two prior studies [71, 105], demonstrating its efficacy for sentiment analysis tasks. After fine-tuning our BERT model (HP-BERT) with the Hinduphobia dataset for Hinduphobic content classification, we will apply multi-stage fine-tuning using the SenWave dataset. Fine-tuning will enable the model ability to perform sentiment analysis on the Global COVID-19 X (Twitter) dataset, refining its understanding of sentiment in the context of Hinduphobic content during the pandemic.
3.1.3 Global COVID-19 X (Twitter) Dataset
@@
The Global COVID-19 X (Twitter) dataset [80] is a comprehensive collection of tweets related to the COVID-19 pandemic, spanning from March 2020 to February 2022. This dataset includes tweets from six countries: Australia, Brazil, India, Indonesia, Japan, and the United Kingdom, as outlined in (Table 4). It captures various public opinions and sentiments about the pandemic, enabling detailed analysis across diverse cultural and regional contexts. This dataset has also been utilised by Chen et al. [17] for sentiment analysis related to Sinophobia (anti-China sentiments) and Chandra et al. [94] for the analysis of antivaccine-related tweets during COVID-19 across different countries. We processed this dataset similarly to the Hinduphobic dataset, ensuring standardised text input essential for the model to accurately capture the contextual nuances in the data. The processed dataset will be used to analyze Hinduphobic tweets across different countries, as well as sentiment analysis in various regions using our HP-BERT model.
@@
Table 4: Tweet counts from six different countries.
@@
3.2 Hinduphobic-BERT Model
BERT [55] is an advanced language model designed to understand the contextual relationships between words, by simultaneously considering both left and right contexts in a sentence. BERT uses a bidirectional attention mechanism based on the Transformer architecture [106], allowing it to read an entire sentence at once and understand the context more effectively. The attention mechanism enables BERT to weigh the importance of different input parts, making it superior for language understanding tasks such as sentiment analysis [107]. The role of BERT in this context is to generate contextualised embeddings for each tweet in the dataset. The pre-training on a massive corpus enables BERT to capture nuanced language features, making it highly suitable for sentiment analysis tasks [94], and detecting hate speech [108] and bias patterns. Furthermore, the BERT model can be easily extended to domain-specific tasks by fine-tuning with specialised datasets.
We develop a Hinduphobic BERT (HP-BERT) model to classify tweets containing Hinduphobic sentiments. The HP-BERT model architecture for detecting Hinduphobic sentiment is shown in (Figure 2). The process begins with the Hinduphobic BERT Tokenizer, which tokenises Hinglish (a blend of Hindi and English), and English tweets into input identifiers (IDs), input type IDs, and input mask. Input IDs represent the tokenised text in numeric form, enabling the model to process it, while input type IDs distinguish between different parts of the input, such as multiple sentences. The attention mask ensures that the model focuses only on meaningful tokens, ignoring padding. These inputs are then passed through a 12-layer encoder to generate contextualised embeddings that are later fed into the fully connected to classify each tweet as either Hinduphobic or Non-Hinduphobic/Neutral. We then use an additional layer for sentiment classification by training (refining) using the SenWave dataset [79]. By combining Hinduphobic detection with sentiment classification, the HP-BERT model enables a comprehensive sentiment analysis of tweets, which also includes the computation of a sentiment polarity score based on predefined weights in (Table 5). This approach allows the model to detect Hinduphobic content while providing detailed sentiment categorisation and sentiment polarity analysis.
@@ 3.2 Hinduphobic-BERT Model
Table 5: Weight ratios for different sentiments for polarity score calculation.
@@
Table 6: Evaluation metrics (mean and standard deviation) for HP-BERT across 30 experimental runs. Metrics include accuracy, F1 score, precision, and recall for various training/testing splits on Hinduphobic COVID-19 X (Twitter) Dataset.
4 Framework
@@
In Stage 1, we obtain three datasets to fine-tune and evaluate our HP-BERT model for sentiment classification, including the i. Hinduphobic COVID-19 X (Twitter) Dataset [103], ii. SenWave Dataset [79], and iii. Global COVID-19 X (Twitter) Dataset [80]. In the case of Hinduphobic COVID-19 X (Twitter) Dataset, we use HITL for manual labelling of 8000 tweets and also develop a model for classification of the leftover tweets.
@@
In Stage 5, we apply our HP-BERT model to perform sentiment analysis on Hinduphobic content, which involves conducting a longitudinal analysis and generating various plots to examine the sentiment distribution and trends across different regions, including sentiment polarity scores. Polarity scores are calculated using the TextBlob library, and custom weight ratios for different sentiment labels as shown in (Table 5) are applied to enhance the analysis. This stage provides deeper insights into how sentiments in Hinduphobic content fluctuate over time and vary geographically.
5 Results
5.1 Technical setup
We use a dropout regularisation rate of $d=0.2$ to prevent overfitting in BERT, obtained from Huggingface [^8] https://huggingface.co/docs/transformers/en/model_doc/bert. The model is trained for 10 epochs with a learning rate of $2\times 10^{-5}$ and a batch size of 8. Additionally, weight decay regularisation with a rate of $\gamma=0.01$ is applied to further mitigate overfitting. We utilise Python-based libraries, including Transformers, NumPy, Pandas, Matplotlib, and PyTorch. The training process is accelerated using an NVIDIA GeForce RTX 2080 Ti GPU (graphics processing unit)with 11264 MiB of memory and an AMD Ryzen 7 5800H processor, along with 16 GB of system memory.
5.1.1 Fine-tuning the HP-BERT Model
We fine-tuned the HP-BERT model for Hinduphobia detection and sentiment analysis using the ”Hinduphobic COVID-19 X (Twitter) Dataset”, which contains 8,000 labelled tweets. The dataset was split into four different training/testing configurations (randomly). The fine-tuned model was evaluated on the test set to assess its performance in classifying Hinduphobic tweets. (Table 6) presents the results from 30 independent experimental runs, reporting the mean and standard deviation for accuracy, F1 score, precision, and recall. The results demonstrate strong performance, with high mean accuracy and balanced F1 scores, indicating the model’s ability to maintain precision and recall. Next, we used the fine-tuned HP-BERT model to predict Hinduphobic tweets in a global dataset.
In (Table 7), we present the results of the fine-tuned HP-BERT model on the SenWave dataset for sentiment analysis, offering additional analysis following Hinduphobic tweet detection. Since this is a multi-label sentiment analysis task, conventional metrics are not applicable. Instead, we report the Label Ranking Average Precision (LRAP) score, F1-score, Jaccard score, and Hamming loss. These metrics have been widely used in earlier studies [94, 71], and our results further validate their findings.

### v2

@@ III-A Datasets
III-A1 Hinduphobic COVID-19 X (Twitter) Dataset
@@
TABLE I: Keywords and Their Association with Negative (Hinduphobic) and Positive & Neutral (Pro-Hindu) Sentiments in Comments
We extracted Hinduphobic tweets and curated a dataset using the X (Twitter) application programmer interface (API) [^2] https://developer.x.com/en/docs/x-api. The dataset spans from April 2020 to January 2024, capturing the context of the COVID-19 pandemic with tweets featuring both Hinglish (a mix of Hindi and English) and English. Each tweet includes terms such as COVID, COVID-19, or coronavirus, establishing a clear relationship between the pandemic and the Hindu community. (Figure 1) illustrates the complete dataset creation and processing pipeline employed in this study.
We employed targeted keywords and hashtags to identify negative (Hinduphobic) and positive/neutral sentiments associated with Hinduism. We present some of the terms identified and their relevance to Hinduphobic sentiments in (Table I), along with positive/neutral terms. The negative terms such as ”Hindu superspreaders” (referring to the Kumbh Mela Hindu festival), ”Hindu rituals causing COVID” (related to temple gatherings), and derogatory references such as ”piss drinker” (associated with cow urine and dung) were used to malign and blame the Hindu community for the spread of COVID-19. Notable hashtags promoting Hinduphobic content included #CoronaJihad, #HindusSpreadCorona, #HinduCOVIDConspiracy, #COVID19HinduHate, and #COVIDHindutva.
We focused on terms associated with Hindu contributions to COVID-19 relief efforts for the positive or neutral sentiment, such as ”Namaste” or ”Namaskar” (a Hindu greeting that avoids physical contact) and references to organisations such as ISKCON (International Society of Krishna Consciousness), RSS (Rashtriya Swayamsevak Sangh), Swaminarayan Temple, and Mahavir Temple. These organizations played a significant role in the relief of COVID-19 by providing food, plasma donations, hospital support, and financial aid. Positive hashtags in this category included #NamasteForSafety, #HindusForHumanity, #ISKCONRelief, #RSSHelpingHands, #TempleAidCOVID, and #HinduCOVIDRelief. We derived the full list of keywords from scholarly sources [96], recent computational studies on religious hate speech detection [97, 98, 99], and validated terminology from peer-reviewed literature on Hinduphobia and religious discrimination [100, 101] with explanations for each term’s association with negative or positive sentiments available in our supplementary materials [^3] https://github.com/pinglainstitute/Hinduphobia-COVID-19-beyond.
Our systematic data collection employed stratified sampling across temporal (early pandemic, Delta variant peak, post-vaccine periods), geographic (six countries with significant Hindu populations: Australia, Brazil, India, Indonesia, Japan, and the United Kingdom to ensure comprehensive regional coverage), and content-type dimensions (news, personal opinions, official statements). We collected approximately 150,000 raw tweets using the Academic Research access with consistent 24-hour collection windows and rate limit compliance (300 requests per 15-minute window) to ensure temporal consistency and data quality across the 44-month collection period.
@@
Hindus are the real reason COVID is spreading! Can’t believe they still gather at temples! #HinduSuperspreaders #COVID19 #Pandemic hindus are the real reason covid is spreading, cannot believe they still gather at temples hindusuperspreaders covid19 pandemic
@@
The thinly veiled Hinduphobia peeping from behind ’secularism’ facade is tiring and oh so cliche. #Secularism #Hinduphobia the thinly veiled hinduphobia peeping from behind secularism facade is tiring and oh so cliche secularism hinduphobia
@@
TABLE II: Examples of tweets related to Hinduphobic before and after preprocessing.
We conducted a comprehensive bias analysis to address potential sampling bias concerns using chi-square goodness of fit tests comparing our tweet distribution against expected distributions based on Hindu population proportions by country, X user penetration rates, and English language proficiency indices. The analysis revealed significant deviations ($\chi^{2}=47.32,p<0.001$), primarily due to India’s overrepresentation; however, this pattern aligns with the country’s large Hindu population (79.8% of 1.4 billion) and high X usage during COVID-19. Temporal bias controls included consistent 24-hour collection windows across all countries and verification that X API rate limits did not systematically affect data collection during high-volume periods. Our statistical analysis showed no significant correlation between collection date and tweet volume bias (Spearman’s $\rho=0.12,p=0.34$), indicating consistent sampling across time periods.
@@
After data collection, we preprocessed the dataset to enhance its quality and relevance by removing irrelevant content, spam, and duplicate entries, along with eliminating web links, hashtags, and profile tags [102, 103, 104]. We also expanded contractions and abbreviations to ensure consistency in data preparation. Additionally, we transformed emojis into text along with other non-text symbols [105, 106] [107]. (Table II) presents examples of tweets before and after preprocessing.
We implemented a sequential annotation workflow designed to leverage both human expertise and model capabilities. The process began with expert human annotators manually labeling 2,000 tweets, establishing a high-quality dataset. We developed annotation guidelines through iterative refinement with cultural context experts to ensure consistent interpretation of Hinduphobic content across different cultural and linguistic contexts. Following the manual annotation phase, we fine-tuned GPT-3.5 Turbo using the 2,000 manually annotated tweets as training data. Our GPT-3.5 Turbo fine-tuning incorporated detailed cultural context, historical background on Hinduphobia, and carefully selected examples representing edge cases and cultural nuances. The training explicitly addressed sarcasm detection, cultural metaphors, and implicit bias recognition to improve automated annotation quality for subsequent labeling tasks.
The fine-tuned GPT-3.5 Turbo model was then deployed to label the remaining tweets from our preprocessed dataset of approximately 15,000 candidates. All GPT-3.5 Turbo-generated labels underwent rigorous human verification, where expert annotators reviewed each prediction and removed incorrectly labeled instances. Through this selective curation process, we retained only 6,000 high-quality GPT-3.5 Turbo labeled tweets that met our annotation standards, ensuring consistency with the manually established gold standard.
Our final annotation quality assessment revealed strong performance metrics: the fine-tuned GPT-3.5 Turbo achieved a 91.7% agreement rate with manual annotation standards. The Cohen’s Kappa between GPT-3.5 labeling and human evaluation was 0.82, indicating substantial agreement. Additional validation involved cross-annotation of 500 tweets by two independent annotators, achieving Cohen’s $\kappa=0.87$, confirming annotation reliability. This comprehensive quality control process achieved 94.2% overall annotation quality using Cohen’s Kappa-based scoring methodology.
Our final annotation quality assessment revealed strong performance metrics: the fine-tuned GPT-3.5 Turbo achieved a 91.7% agreement rate with manual annotation standards. The Cohen’s Kappa between GPT labeling and human evaluation was 0.82, indicating substantial agreement. Additional validation involved cross-annotation of 500 tweets by two independent annotators, achieving Cohen’s $\kappa=0.87$, confirming annotation reliability. This comprehensive quality control process achieved 94.2% overall annotation quality using Cohen’s Kappa-based scoring methodology. (Table III) provides a comprehensive overview of our multi-phase annotation workflow, detailing the progression from initial manual annotation through AI-assisted labeling to final quality validation.
TABLE III: Dataset Annotation Summary and Quality Metrics
Phase Tweets Method Agreement Quality
Manual 2,000 Expert Human - 100%
GPT Fine-tuning 2,000 Training Data - -
GPT Labeling 15,000 Fine-tuned Model 91.7% -
Human Verification 6,000 Quality Selection $\kappa$ = 0.82 94.2%
Cross-Validation 500 Dual Annotators $\kappa$ = 0.87 -
Final Dataset 8,000 Hybrid Approach $\kappa$ = 0.87 94.2%
Figure 2: HP-BERT model architecture for detecting Hinduphobic sentiment, featuring tokenisation (encoding) of Hinglish/English X using a set of encoders. It includes a fully connected layer for classifying X as Hinduphobic or Non-Hinduphobic/Neutral, and the final layer provides sentiment classification into different categories.
@@
The ”Hinduphobic COVID-19 X (Twitter) Dataset” is publicly available on Kaggle under a Creative Commons license. To ensure user privacy and platform compliance, the dataset contains only tweet IDs and classification labels, with all personally identifiable information (usernames, profile details, location data) removed. The final dataset contains 8,000 tweets (4,600 labeled as Hinduphobic and 3,400 as positive/neutral), published on Kaggle  [^4] https://www.kaggle.com/datasets/ashutoshsingh22102/hinduphobic-covid-19-x-twitter-dataset-india. This rigorous annotation process was necessary due to the challenge of accurately identifying hate speech, which is inherently language and culture-dependent. Existing models, such as HateBERT [76] and COVID-HateBERT [108], may struggle to detect nuanced expressions of Hinduphobic content with hidden meanings. Correctly identifying instances of hate speech is essential to maintaining data integrity, ensuring accurate interpretation, and minimizing the risk of inaccuracies. This dataset serves as the foundation for fine-tuning the HP-BERT model, specifically designed to improve the detection of Hinduphobic content.
III-A2 SenWave dataset
@@
The SenWave dataset [57] contains over 104 million tweets and Weibo messages related to COVID-19 in six different languages: English, Spanish, French, Arabic, Italian, and Chinese. The data was collected between March 1 and May 15, 2020, to analyse social media posts surrounding the pandemic. It includes 10,000 English and 10,000 Arabic tweets labelled into ten sentiment categories: optimistic, thankful, empathetic, pessimistic, anxious, sad, annoyed, denial, official report, and joking. This comprehensive labelling facilitates in-depth analysis of the diverse emotional responses exhibited during the pandemic. The effectiveness of the SenWave dataset has been validated in two prior studies [82, 109], demonstrating its efficacy for sentiment analysis tasks. After fine-tuning our BERT model (HP-BERT) with the Hinduphobia dataset for Hinduphobic content classification, we will apply multi-stage fine-tuning using the SenWave dataset. Fine-tuning will enable the model to perform sentiment analysis on the Global COVID-19 X (Twitter) dataset, refining its understanding of sentiment in the context of Hinduphobic content during the pandemic.
III-A3 Global COVID-19 X (Twitter) Dataset
@@
The Global COVID-19 X (Twitter) dataset [62] is a comprehensive collection of tweets related to the COVID-19 pandemic, spanning from March 2020 to February 2022. This dataset includes tweets from six countries: Australia, Brazil, India, Indonesia, Japan, and the United Kingdom, as outlined in (Table IV). It captures various public opinions and sentiments about the pandemic, enabling detailed analysis across diverse cultural and regional contexts. This dataset has also been utilised by Chen et al. [17] for sentiment analysis related to Sinophobia (anti-China sentiments) and Chandra et al. [95] for the analysis of antivaccine-related tweets during COVID-19 across different countries. We processed this data set similarly to the Hinduphobic data set, ensuring standardized text input essential for the model to accurately capture contextual nuances in the data. The processed dataset will be used to analyze Hinduphobic tweets across different countries, as well as sentiment analysis in various regions using our HP-BERT model.
@@
TABLE IV: Tweet counts from six different countries.
@@
III-B Hinduphobic-BERT Model
BERT [66] is an advanced language model designed to understand the contextual relationships between words by simultaneously considering both left and right contexts in a sentence. BERT uses a bidirectional attention mechanism based on the Transformer architecture [110], allowing it to read an entire sentence at once and understand the context more effectively. The attention mechanism enables BERT to weigh the importance of different input parts, making it superior for language understanding tasks such as sentiment analysis [111]. The role of BERT in this context is to generate contextualised embeddings for each tweet in the dataset. The pre-training on a massive corpus enables BERT to capture nuanced language features, making it highly suitable for sentiment analysis tasks [95], and detecting hate speech [112] and bias patterns. Furthermore, the BERT model can be easily extended to domain-specific tasks by fine-tuning with specialised datasets.
We develop a Hinduphobic BERT (HP-BERT) model to classify tweets containing Hinduphobic sentiments. The HP-BERT model architecture for detecting Hinduphobic sentiment is shown in (Figure 2). The process begins with the Hinduphobic BERT Tokenizer, which tokenises Hinglish (a blend of Hindi and English) and English tweets into input identifiers (IDs), input type IDs, and input mask. Input IDs represent the tokenised text in numeric form, enabling the model to process it, while input type IDs distinguish between different parts of the input, such as multiple sentences. The attention mask ensures that the model focuses only on meaningful tokens, ignoring padding. These inputs are then passed through a 12-layer encoder to generate contextualized embeddings that are later fed into the fully connected layer to classify each tweet as either Hinduphobic or Non-Hinduphobic/Neutral. We then use an additional layer for sentiment classification by training (refining) using the SenWave dataset [57]. The HP-BERT model enables a comprehensive sentiment analysis of tweets by combining Hinduphobic detection with sentiment classification, which also includes the computation of a sentiment polarity score based on predefined weights in (Table V). This approach allows the model to detect Hinduphobic content while providing detailed sentiment categorization and sentiment polarity analysis.
@@ III-B Hinduphobic-BERT Model
TABLE V: Weight ratios for different sentiments for polarity score calculation.
@@
TABLE VI: Evaluation metrics (mean and standard deviation) for HP-BERT across 30 experimental runs. Metrics include accuracy, F1 score, precision, and recall for various training/testing splits on Hinduphobic COVID-19 X (Twitter) Dataset.
IV Framework
@@
In Stage 1, we obtain three datasets to fine-tune and evaluate our HP-BERT model for sentiment classification, including the i. Hinduphobic COVID-19 X (Twitter) Dataset [113], ii. SenWave Dataset [57], and iii. Global COVID-19 X (Twitter) Dataset [62]. In the case of the ”Hinduphobic COVID-19 X (Twitter) Dataset”, we use HITL for manual labelling of 8000 tweets and also develop a model for classification of the leftover tweets.
@@
In Stage 5, we apply our HP-BERT model to perform sentiment analysis on Hinduphobic content, which involves conducting a longitudinal analysis and generating various plots to examine the sentiment distribution and trends across different regions, including sentiment polarity scores. Polarity scores are calculated using the TextBlob library, and custom weight ratios for different sentiment labels, as shown in (Table V) are applied to enhance the analysis. This stage provides deeper insights into how sentiments in Hinduphobic content fluctuate over time and vary geographically.
V Results
V-A Technical setup
We use a dropout regularization rate of $d=0.2$ to prevent overfitting in BERT, obtained from Huggingface [^5] https://huggingface.co/docs/transformers/en/model_doc/bert. The model is trained for 10 epochs with a learning rate of $2\times 10^{-5}$ and a batch size of 8. Additionally, weight decay regularization with a rate of $\gamma=0.01$ is applied to further mitigate overfitting. We utilize Python-based libraries, including Transformers, NumPy, Pandas, Matplotlib, and PyTorch. The training process is accelerated using an NVIDIA GeForce RTX 2080 Ti GPU (graphics processing unit)with 11264 MiB of memory and an AMD Ryzen 7 5800H processor, along with 16 GB of system memory.
V-A1 Fine-tuning the HP-BERT Model
We fine-tuned the HP-BERT model for Hinduphobia detection and sentiment analysis using the ”Hinduphobic COVID-19 X (Twitter) Dataset”, which contains 8,000 labeled tweets. The dataset was split into four different training/testing configurations (randomly). The fine-tuned model was evaluated on the test set to assess its performance in classifying Hinduphobic tweets. (Table VI) presents the results from 30 independent experimental runs, reporting the mean and standard deviation for accuracy, F1 score, precision, and recall. The results demonstrate strong performance, with high mean accuracy and balanced F1 scores, indicating the model’s ability to maintain precision and recall. Next, we used the fine-tuned HP-BERT model to predict Hinduphobic tweets in a global dataset.
In (Table VII), we present the results of the fine-tuned HP-BERT model on the SenWave dataset for sentiment analysis, offering additional analysis following Hinduphobic tweet detection. Since this is a multi-label sentiment analysis task, conventional metrics are not applicable. Instead, we report the Label Ranking Average Precision (LRAP) score, F1-score, Jaccard score, and Hamming loss. These metrics have been widely used in earlier studies [95, 82], and our results further validate their findings.